OpenStudy (anonymous):

How do you answer y=sin(x) for angle -90, -60, and -45 degrees?

OpenStudy (debbieg):

I guess you mean: sin(-90)= ? sin(-60)= ? sin(-45)= ? It kind of depends on what methods you have covered. If your teacher is stressing the unit circle, then those are on there, so you would look at the y-coordinate of each point of intersection with the circle. Alternatively, for -90 you have a quadrantal angle. since \(\sin\alpha=y/r\) then you have to notice that, for any point on the negative side of the y-axis, y/r=-1. For the other two, those are not quadrantal so I would use their reference angles to find the function values. Sine is negative in Q4, so you have: sin(-60)=-sin(60) And sin(-45)=-sin(45) For sin(60) and sin(45) then, you could use SOHCAHTOA with the appropriate "special triangle":